OpenStudy (anonymous):

how do i solve the square root of x with the square root of x and index of 3 in the inside?

OpenStudy (chaise):

\[\sqrt{x \times \sqrt[3]{x}}\] is this right?

OpenStudy (anonymous):

yes but it doesnt have the multipliction (x) in the problem. if that makes sense. sorry i dont know how to make a sqaure root sign.

OpenStudy (chaise):

It doesn't need a multiplication sign, because they are side by side it is assumed multiplication. \[\sqrt{x \times \sqrt[3]{x}}\] \[=\sqrt{x \times x^{\frac{1}{3}}}\] \[=\sqrt{x^{\frac{4}3}}\] \[=x^{\frac{2}{3}}\] \[\sqrt[3]{x^2}\]

OpenStudy (anonymous):

if you dont mind explaining how you got 2/3 i understood all the way up to there. ?